The radius of the circle having centre at (2,1) whose one chord is the diameter of the circle x2 +y2 -2x-6y+6=0

Grade:12th pass

1 Answers

Vikas TU
6 years ago
X2+y^2 -2x-6y+6=0
Rearranging,
(x^2-2x+1)+(y^2-6y+9)-4=0
Forming a perfect square,
or, (x-1)^2+(y-3)^2=22
thus the radius of the smaller given circle is 2 and its centre is (1,3).
distance b/w the two centres = (1+4)^1/2 = 5^1/2
thus the radius of required circle = (5 + 4)^1/2 = 3.